How would you priduce a 10^-1 dilution of a 3 mL bacterial sample using the entire 3mL volume?
Answer #1

Dilution = initial volume / final volume

Final volume = initial volume + volume of diluent

Dilution = 10^-1

Initial volume = 3 mL

10^-1 = 3 / final volume

Final volume = 3 / 10^-1 = 30 mL

30 = 3 + volume of diluent

Volume of diluent = 30 - 3 = 27 mL

So, to 3 mL of sample, add 27 mL of diluent.

Please rate high.